Mullet Technologies is considering whether or not to refund a $125 million, 12% coupon, 30-year bond issue that was sold 5 years ago. It is amortizing $6 million of flotation costs on the 12% bonds over the issue's 30-year life. Mullet's investment banks have indicated that the company could sell a new 25-year issue at an interest rate of 9% in today's market. Neither they nor Mullet's management anticipate that interest rates will fall below 9% any time soon, but there is a chance that rates will increase. A call premium of 15% would be required to retire the old bonds, and flotation costs on the new issue would amount to $6 million. Mullet's marginal federal-plus-state tax rate is 40%. The new bonds would be issued 1 month before the old bonds are called, with the proceeds being invested in short-term government securities returning 6% annually during the interim period. Conduct a complete bond refunding analysis. What is the bond refunding's NPV? Do not round intermediate calculations. Round your answer to the nearest cent. $

What factors would influence Mullet's decision to refund now rather than later?
Answer #1

In order to conduct a complete bond refunding analysis, we will have to compute the following three components. Let's calculate them one by one:

  1. Total investment outlay
  2. Annual cash flows comprising of:
    1. Net amortization tax effects
    2. Net post tax interest savings
  3. Calculate NPV of bond refunding using total investment outlay and annual cash flows

Part 1: Total initial investment outlay

  • Pre tax call premium to be paid on refund = Call premium rate x face value of old issue = 15% x $ 125,000,000 = $18,750,000.00
  • Post tax call premium to be paid = Pre tax call premium to be paid x (1 - tax rate) = $18,750,000 x (1 - 40%) = $11,250,000.00
  • Floatation cost on new issue = $6,000,000.00
  • Balance old floatation cost = (Old floatation cost / Life of the old bond) x balance life = 6,000,000 / 30 x (30 - 5) = $5,000,000.00
  • Since we are retiring the old bonds, we will get the tax benefit by expensing the balance floatation cost immediately. Tax benefit on expensing the balance floatation cost = Balance floatation cost x tax rate = 5,000,000 x 40% = 2,000,000
  • Since there is a time gap of 1 month between the new bond issue and retiral of old bonds, we will have to continue to pay the interest on the old bond issue for this one month. However this interest will also give us an interest tax shield. Hence, post tax interest on old bonds for 1 month = Face Value x interest rate of old bond x 1 / 12 x (1 - Tax rate) = 125,000,000 x 12% x 1 / 12 x (1 - 40%) = 750,000
  • New bonds issued will be invested in short-term government securities returning 6% annually during the interim period of 1 month. However this interest income will be subjected to tax as well. Hence, post tax interest income = Face value of new bond issue x short term government securities interest rate x 1 / 12 x (1 - Tax rate) = 125,000,000 x 6% x 1/12 x (1 - 40%) = 375,000

Thus, total initial investment outlay = Post tax call premium paid + New floatation cost - Tax savings by expensing the balance floatation costs on old bonds + post tax interest paid on old bonds in the interim period of 1 month - post tax interest earned on proceeds from new bond issue invested in short-term government securities during the interim period of 1 month = 11,250,000 + 6,000,000 - 2,000,000 + 750,000 - 375,000 = $15,625,000.00. Please note that this is an outlay i.e. a cash outflow .

Part 2: Annual cash flows. This comprises of two sub parts:

  • Total amortization tax effects
    • Annual floatation cost on new issue = Total floatation cost of new issue / Life of new issue = 6,000,000 / 25 = 240,000
    • Annual floatation cost foregone on old issue = Total floatation cost of old issue / Life of old issue = 6,000,000 / 30 = 200,000
    • Incremental annual floatation cost that will be amortised = 240,000 - 200,000 = 40,000
    • Total amortization tax effect = Tax saved due to incremental annual floatation cost amortisation= 40,000 x Tax rate = 40,000 x 40% = 16,000
  • Figure out net post tax interest savings
    • Annual interest on old bonds = 125,000,000 x 12% = 15,000,000
    • Annual interest on new bonds = 125,000,000 x 9% = 11,250,000
    • Annual interest saved = 15,000,000 - 11,250,000 = 3,750,000
    • Net Post tax interest saving = Pre tax interest saving x (1 - Tax rate) = 3,750,000 x (1 - 40%) = 2,250,000

Thus annual cash flow = total amortization tax effects + Net Post tax interest saving = 16,000 + 2,250,000 = $ 2,266,000.00

Part 3: We are now ready to calculate NPV

Initial investment outlay (as calculated in part 1 above) = $ 15,625,000.00

Annual post tax cash inflows = $ 2,266,000.00 (as calculated in part 2 above)

NPV = - Initial investment + PV of all the future annual post tax cash inflows

For PV of all the future annual post tax cash inflows:

Discount Rate = short-term government securities rate = 6%

Period = 25 years

Payment = 2,266,000

Use excel function "PV" to calculate the PV of all the future annual post tax cash inflows = PV (Rate, Period, Payment, FV) = PV(6%, 25, -2266000,0) =   $ 28,967,085.05

Hence, NPV = - Initial investment + PV of all the future annual post tax cash inflows = -15,625,000 + 28,967,085.05 =   $ 13,342,085.05

Hence, the bond refunding's NPV = $ 13,342,085.05

What factors would influence Mullet's decision to refund now rather than later?

Every variable that has played a role in the calculation above would influence the decision to refund now rather than later:

  • Call premium rate
  • Balance floatation cost on the old bond
  • Floatation cost on the new bond issue
  • Coupon rate of new bond in comparison to the coupon rate of old bond issue
  • Return on short-term government securities